OpenStudy (anonymous):

what is the midpoint of a segment whose endpoints are (13,-24) and (17,-6)?

OpenStudy (anonymous):

...i need helpp

OpenStudy (anonymous):

;ife saver

OpenStudy (tanner23456):

We use the same equation as last time x coordinate = (13+17)/2 y coordinate = (-24+-6)/2

OpenStudy (anonymous):

15 15

OpenStudy (phi):

close, but I think the y values should be negative

OpenStudy (phi):

notice the average of the y values is (-24 + -6)/2 = -30/2 = -15